FEL Upgrade Project Weekly Brief

February 25-March 1, 2002




The procurement package for the optical cavities for the 10 kW Upgrade were sent out this week.


We completed revisions of the our primary contractual agreement between the Navy and DOE for the FEL activities at Jefferson Lab (the "MOA"), that would fund additional work in FY02 and FY03 including:

(1) commissioning of the FEL Upgrade

(2) design and construction of a short wavelength (1-2 micron) wiggler

(3) completion of the 3rd cryomodule for the FEL linac

(4) upgraded drive laser for the photocathode gun (JTO funded)

(5) design and testing of cryogenically cooled optical cavity mirrors (JTO funded)

The revised MOA includes work statements and cost summaries for the above activities. These documents have been forwarded to our DOE and ONR contracting officers for approval.

This is the last week of the 2002 session of the Virginia General Assembly, so we had to respond to several requests for information concerning our FEL activities that are funded by the Commonwealth of Virginia.

FEL project management and AES,Inc have agreed upon the joint work statement for a CRADA on 100 mA injector development. The document goes on for the final packaging needed for the governmental (DOE/DOD) reviews. A support letter enumerating the JLab tasks was forwarded to AES.

WBS 3 (Beam Physics):

Consideration of endloop dipole (GX, GQ, GY) performance continues. All appear to have adequate field uniformity over the full operating range (80-210 MeV/c). All appear, however, to

run somewhat saturated at full (210 MeV/c) excitation. As field and field integral matching was done at this operating point, all fail to stay within tracking specification (1 ppt) for field and field integral as excitations are lowered into the 80-145 MeV range.

Initial analysis suggests that GX & GQ magnets can be well matched in field integral for IR operation over 80-190 MeV/c by appropriate reduction of the steel length (about 3 mm for the GX and 5 mm for the GQ). This leads to very good field performance in the GQ (less than 1 ppt from design), while leaving the GX a factor of 2 or so out of spec in field. This is likely acceptable - the GX is at a location with modest beam envelopes and low dispersion; the focussing errors associated with deviations in field will not strongly influence machine performance. With "shaved" lengths at 210 MeV/c, field integrals in the GX and GQ fail to track at the 1/4% level, and core fields fail to track at the 1/3% level. Should this prove operationally unacceptable (either through excessive correction requirements or excessive focussing errors), shimming the GX with a single 2.5 mm shim (on the IR path, at the exit of the magnet) and shimming the GQ with a pair of 2.5 mm shims will bring both back within specification.

UV operation of the GX is also influenced by saturation. Core field matching at full excitation (210 MeV/c) will require ~2% offset of the bypass GX from the buss current; this falls to ~0.1% at 145 MeV/c. In either case, the magnet also fails to track in field integral by an amount consistent with ~3 mm length error. Should the aforementioned "shave" be done, this will be compensated at all excitations. Core field matching can in principle be accomplished through the use of a shunt. In practice, it may be less risky to energize the "shorted" coil using a trim supply, inasmuch as the required correction (0.1%) at low energy suggests that the shunt may run out of range and a bipolar trim be required if the magnetic model proves even slightly inaccurate.

GY performance poses a similar challenge - nominally designed to shunt 1.5 % of the current at full excitation, the required GY shunt falls to order 0.1% at 80 to 145 MeV/c. In practice, the magnet may run outside the available shunt range. This may be addressed two ways - a correction coil winding on the yoke and/or a small reduction in gap. Given that a single "backleg" turn will require 7m of wire, the latter solution may be preferable.

GW performance will be similarly examined to ensure it matches adequately with all this. George B. points out that not only will all this likely produce a properly functioning machine, it gives the WBS 3 manager magnets that are just far enough from spec that he can whine about them indefinitely, even though they'll work just fine.

WBS 4 (Injector):

The gun chamber has had its flanges welded on. The port screens have been welded, shaped and are being fitted up. The cathode support tube is polished, cleaned and ready for implantation.

A GaAs wafer was hydrogen cleaned and afterwards studied under the SEM. No signs of dust were observed, although some "flake like" features appear scattered all over the surface. No emitters were found by the DC Field Emission Scan at 35 MV/m, except for a small emitter on the edge of the wafer. These preliminary results are encouraging since the electric field in the upgrade gun will be 6-10 MV/m. A new Scan is in progress at 60 MV/m.

Gun HVPS - The design of the combined Running & Conditioning Resistor mechanism is nearly finished. A design review will be held next week. The breakdown voltage and expected safety margin at several locations has been calculated. The bids for the new HVPS Tank closed this week and a contract was placed with Industrial Alloy Fabrication located in Richmond. It is due May 10th.




WBS 6 (RF):

Injector HVPS - The Electricians finished wiring the high current primary side of the HVPS. New wiring was pulled and wired to the build's switchgear. For safety reasons, the final connection to the HVPS will not be made until it is needed to checkout the HVPS. Design of the PLC system and code is progressing.

Injector Klystrons - Brackets were made and used to install the new Circulators. Discussions are ongoing on how to install the larger LCW taps. New 8" LCW piping is being considered for the first 30' in the RF Gallery. The removal of an orifice to measure LCW flow is being considered also. These changes are to provide more flow for the upgrade equipment.

WBS 8 (Instrumentation):

Nine additional beamviewers have been delivered for assembly onto girders bringing the total to fourteen. Beam Position Monitors will be tested and delivered for assembly early next week.

The Beamloss Monitoring System design review is in progress. Procurement and fabrication is expected to begin soon. We are also moving into the design phase of an upgrade to the Analog Monitoring System. Work on a system drawing for the Laser Safety System has resumed as well.

Trim magnet cable mapping is well underway. We expect to begin pulling new cables and re-routing old ones as soon as the underway portion of the mapping of the floor in the vault is complete.

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

o At Magnet Enterprises International in Oakland CA. the second coil will be potted on Monday.

o Return legs remain being ground to their final size. Top and bottom slabs will finish rough machining early next week. Field clamp material is in and those parts are being machined.

Injector Dipoles (DU/DV)

o No news from Wang NMR on coils.

o Master Machine will pick up material from us on Monday.

o DULY Research completed the Revision A of the drawings for the Small Injector Dipole (GV) to give the final adjustment to meet specifications. We are checking the changes.

Arc Dipoles (GY, GX, GQ)

o Electronic procurement of Bend (GX) and Reverse Bend (GQ) Dipoles continued. We have given vendors an additional week to respond because two vendors requested the extension.

o David Douglas has has almost completed his analysis of the use of the Bend and Reverse Bend Dipoles (GX & GQ) at all levels of excitation after magnetic model analysis from AES indicated the magnets remain linear to that point. His conclusion is that the faces may need to be shaved by about 2 mm with shim back only necessary for use at 210 MeV/c.

o Wang NMR of CA is now starting to make the GY coil winding mandrel.

o Bosma Machine is preparing for rough machining of the core pieces. Shim pieces are cut to size.



3 inch quad (QX)

o Eight of the QX magnet were measured. All met the field quality requirements. However, they are a factor of 4 out of specification in matching from magnet to magnet that would allow one excitation curve to characterize them all. They are able to be divided in groups with individual excitation curves and keep track of those in the software.

Trim Quad (QT)

o Milhous Control of Virginia started winding the first magnet at last. The custom magnet wire we supplied was out of specification. New wire that we supplied restarted the process. We are working with the wire manufacturer so that we can return the out of spec wire for credit.

Sextupole (SF)

o DULY Research has not advanced this week because their analysis expert was back in Russia tending to his wife and new born daughter.

Octupole (OT)

o Work on this magnet is on furlough until the sextupole is designed.

Beam Line and Vacuum

o AES continues drawing up the chambers throughout the arcs. The task is about 60% complete.

o Design of the Arc Dipole Stands continued.

o The Alignment Group established their expanded network in the FEL Enclosure. They will lay down the position of stands during short intervals over the next month.

WBS 11 (Optics):

The drawings for the optical cavity assemblies were released for procurement this week. We continue work on the insertable mirror drawings. We are awaiting quotes on some of the mirrors for the OCMMS before final release of the drawings. The other long lead item (the calcium fluoride viewports) were ordered. We received the pillow blocks, another major component for the optical cavity, this week. These will be cleaned for UHV service next week. We received the modified ultraviewer components; and discovered the brackets had been bead blasted. After discussions with various vacuum-savvy folks, we have decided to have them remade.

We held more meetings on collimator designs, LCW requirements in the vault and upstairs, and electrical schematics for the optical cavity assemblies.

The pump laser for the mirror test stand (MTS) and the optical table supporting the laser interferometer and the MTS have improved vibration-isolation. This, and additional stiffening of the MTS floor have greatly lowered the amount of noise we see on the interferograms. Checks of the laser interferometer after the service done last week indicate that it is fixed.

Other Activities:

Several of the group attended one-day training for the new program management software the division wishes us to use.